# Exploring Spatial Representation to Enhance LLM Reasoning in Aerial Vision-Language Navigation

### v2

@@ Abstract
Aerial Vision-and-Language Navigation (VLN) is a novel task enabling Unmanned Aerial Vehicles (UAVs) to navigate in outdoor environments through natural language instructions and visual cues. However, it remains challenging due to the complex spatial relationships in outdoor aerial scenes. In this paper, we propose a training-free, zero-shot framework for aerial VLN tasks, where the large language model (LLM) is leveraged as the agent for action prediction. Specifically, we develop a novel Semantic-Topo-Metric Representation (STMR) to enhance the spatial reasoning capabilities of LLMs. This is achieved by extracting and projecting instruction-related semantic masks of landmarks onto a top-down map, which contains location information about surrounding landmarks. The map is then transformed into a matrix representation with distance metrics, serving as the text prompt to the LLM for action prediction in response to the given instruction. Experiments conducted in real and simulation environments have successfully proved the effectiveness and robustness of our method, achieving at most 10.8% improvements (absolute) in Oracle Success Rate (OSR) on AerialVLN-S and Openfly dataset.
I Introduction
@@
The Aerial Vision-and-Language Navigation (Aerial VLN) [liu_2023_AerialVLN] emerges as a groundbreaking task. It enables unmanned aerial vehicles (UAVs) to interpret natural language instructions and corresponding visual information to navigate in outdoor environments. This technology can eliminate the necessity for manual UAV operation by human pilots, clearly mitigating the barriers to human-UAV interaction and potentially benefitting search, rescue, and delivery tasks.
@@
Recently, VLN tasks have been well-developed. Considering the powerful reasoning capability of large language models (LLMs) [ahn2022can, song2023llm, driess2023palm, huang2022inner], several VLN methods have started to use LLMs as agents to parse instructions and predict actions [zhou2024navgpt, shah2023lm, li2024tina, lin2024navcot]. Specifically, existing methods attempt to describe the visual observations with text to enhance the LLM’s scene understanding ability. NavGPT [zhou2024navgpt] converted visual scene semantics into text description for the LLM to conduct high-level planning. MapGPT [chen2024mapgpt] introduced a topological graph described in language as text prompts to an LLM, guiding low-level planners to explore new nodes of a predefined graph of the environment.
Figure 1: The pipeline to obtain STMR. Observation image, semantic mask, and depth image (a) can be aggregated in the top-down map (red dashed box) (b), which captures the Semantic-Topo-Metric information of the neighborhood of the current position, further generating a 20x20 matrix representation (c) for LLM reasoning.
@@
Although existing VLN methods have made significant progress, most of them are designed for indoor or ground-based outdoor environments and pay less attention to aerial VLN tasks. Practically, the rich semantics, complex spatial relationships, and the large-scale nature of outdoor 3D spaces make it challenging for existing VLN methods to adapt to aerial scenarios. As shown in  Fig. 1 (a), the aerial scene can be quite complex, which may lead to incorrect reasoning for several reasons, such as an overemphasis on instruction-irrelevant objects, or a failure to capture the contextual relationships between different areas. Furthermore, the vast scale of aerial landscapes can further complicate navigation, making local-perspective-based VLN methods unable to comprehend the environment. Thus, the development of VLN methods that integrate both effective semantic information and precise spatial representations is urgently needed to enhance the adaptability and accuracy in aerial scenarios.
@@
To overcome these challenges, we propose a zero-shot LLM-based aerial VLN framework, which takes natural language instructions, RGB images, and depth images as input, and directly makes action predictions (e.g., forward 10 meters) through single-step reasoning updates. Specifically, a Semantic-Topo-Metric Representation (STMR) is designed for LLMs spatial reasoning. Leveraging the developments in open-vocabulary perception models [zhang2018review, jiang2018deep, minderer2022simple, cheng2024yolo], we first extract instruction-related landmarks and obtain corresponding semantic masks by perception models, i.e., Grounding Dino [liu2023grounding] and Tokenize Anything [pan2023tap]. After that, semantic masks obtained from each step are gradually projected into a top-down map using depth and pose transformation as shown in Fig. 1 (b). This projected map encompasses both the UAV’s travel trajectory and spatial information, i.e., locations of surrounding objects. To integrate rich visual information and topology into the text prompt while maintaining its simplicity, we separate the top-down map centered at the UAV’s current position into grids and substitute each grid with a semantic number. This process transforms the top-down map into a matrix representation, tailored to serve as a spatial prompt for the LLM. As can be seen from Fig. 1 (c), the matrix representation encompasses topological, semantic, and metric information, and is input to an LLM together with historical actions, sub-goal instructions, and other information to infer the next action. Compared with other learning-based and LLM-based methods, the proposed method effectively improves the accuracy of spatial reasoning.
@@
  • To our knowledge, we design the first LLM-based training-free framework for the aerial VLN task, facilitating the development of UAV navigators. Without extra action planner, the proposed framework allows for easy integration.
@@
II Related Work
Vision-Language Navigation (VLN)
@@ II Related Work
VLN aims to enable autonomous agents to navigate complex environments by understanding and executing natural language instructions based on visual context. Early VLN methods use sequence-to-sequence LSTMs to predict low-level actions [anderson2018vision] or high-level actions from panoramas [fried2018speaker]. Several attention processes have been proposed [qi2020object, hong2020language, an2021neighbor] to enhance the process of learning visual textual correspondence. Reinforcement learning is also explored to improve policy learning [wang2018look, tan2019learning, wang2020soft]. Besides, transformer-based architecture have shown superior performance to long-distance contextual information [hao2020towards, majumdar2020improving]. More recent works [zhou2024navgpt, chen2024mapgpt] leverage the reasoning and dialogue capabilities of LLMs, achieving great progress. However, most of them are developed under a ground-based, discrete VLN setting, which simplifies navigation by traversing on a predefined graph on the ground. This limits the UAVs’ free movement space in the real world. Aerial VLN remains challenging due to the large-scale and complex environments.
UAV navigation
@@
Unmanned Aerial Vehicle (UAV) navigation has seen a surge of interest over the past few decades. Guided with instructions, Blukis et al. [blukis2018following] use imitation learning for low-level velocity command prediction. [blukis2018mapping] predicts interpretable position-visitation distributions to guide UAV’s control actions, using a combination of supervised and imitation learning for efficient training. Misra et al. [misra2018mapping] decompose instruction execution in two stages, separately using supervised learning for goal prediction and policy gradient for action generation. AerialVLN [liu_2023_AerialVLN] contributes a much more challenging aerial VLN dataset focusing on outdoor aerial environments, and provides a look-ahead guidance method as the baseline. OpenFly [openfly] constructs a large-scale dataset across several high-quality scenes, and proposes a keyframe-aware VLN model. Despite the progress, the generalizability and performance of these methods still require improvement.
LLMs for robot planning and interaction
@@
Most recently, LLMs have demonstrated impressive capabilities in understanding and reasoning. To leverage these capabilities, several promising methods have been proposed for applying LLMs in robotic systems. A few methods involve using LLM-generated rewards optimized in simulation to improve control [huang2023voxposer, yu2023language]. Others utilize LLM-selected subgoals as an abstraction to enhance policies for navigation [dorbala2022clip, chen2023open] and manipulation [cui2022can, shridhar2022cliport]. Additionally, research has explored the use of LLMs to generate executable code for control and perception primitives [singh2023progprompt, wu2023tidybot, liang2023code]. Despite their potential, LLMs are still prone to confidently hallucinating outputs, such as referring to objects not observed in the scene [zeng2022socratic]. In order to alleviate the hallucinating phenomenon, we propose a matrix-based representation containing topological, semantic, and metric information for better prompting the LLMs.
III Method
@@
In this paper, we introduce a novel zero-shot framework that leverages Large Language Models (LLMs) for action prediction in aerial VLN tasks. As shown in Fig. 2, our framework mainly consists of three modules. The sub-goal extracting module decomposes language instructions into several sub-goals, facilitating step-by-step reasoning and navigation. The Semantic-Topo-Metric Representation (STMR) module represents the outdoor environment as a matrix containing semantic, topological, and metric information for the prompt to an LLM. Finally, the LLM planner outputs its current thoughts and actions by using the prompts from the aforementioned two modules along with the task description and history as inputs.
III-A Problem Formulation
@@
In an aerial VLN task, the problem is formulated as a free-form language instruction guided navigation. At the beginning of each episode, the UAV is placed in an initial pose $P=[x,y,z,\phi,\theta,\psi]$, where $(x,y,z)$ denotes the UAV’s position and $(\phi,\theta,\psi)$ represents pitch, roll, and yaw of the UAV’s orientation. A natural language instruction $L$ is provided to specify the path that a UAV should follow. To achieve the navigation goal, the UAV considers both the instruction and visual perceptions, and predicts an action from the action space (right, left, up, down, forward, backward) with corresponding parameters at each time step $t$. Navigation ends when the UAV predicts a ‘Stop’ action or reaches a pre-defined maximum action number. Following [liu_2023_AerialVLN], the navigation is recognized as a success if the UAV stops at a location that is less than 20 meters from the target location.
Figure 2: Our method consists of three modules, i.e. sub-goal extracting, Semantic-Topo-Metric Representation, and LLM planner. First, we prompt an LLM to break down long instructions and extract specific landmarks, which will be classified by perception models, generating corresponding semantic masks. Then we project these masks into a top-down map using the UAV-view pose and depth information. The map is further compressed into a matrix representation. Finally, we prompt LLM with well-designed instructions to enable effective reasoning.
III-B Semantic-Topo-Metric Representation (STMR)
@@
Previous LLM-based VLN methods use natural language to describe current observations, or nodes and edges of a topological graph to model the spatial information of the environment. However, in open scenarios, simple directional words such as “next to” or “aside” are not enough to describe complex spatial relationships, which can easily introduce ambiguity into LLMs. To address this challenge, we introduce the STMR to enhance the spatial-aware reasoning capability of LLMs. Specifically, STMR incrementally takes an RGB image $I_{t}^{R}$ and a depth map $I_{t}^{D}$ as input from each step, and generates a dynamically updated matrix representation with semantic, topological, and metric information as its output. The details of STMR are presented as follows.
III-B1 2D Visual Perception
@@
$\{m^{(t)},\ h^{(t)}\}=\operatorname{TA}\left({I_{t}^{R}},\ \operatorname{GD}\left({I_{t}^{R}},\ \operatorname{LE}({L})\right)\right).\$ (1)
To improve the simplicity of the semantic masks and reduce the misleading of LLM reasoning caused by numerous open-vocabulary categories, we propose a text-matching method to mitigate over-classification. As shown in Fig. 3, our method involves vectorizing the landmarks extracted from the instruction and the captions $h^{(t)}$ generated in each $I_{t}^{R}$. Then, we calculate the cosine similarity between these vectorized landmarks and captions using TF-IDF [qaiser2018text]. If the similarity score exceeds the threshold $\tau$ (0.8), the landmark is classified as visible in the current view, and other irrelevant masks will be ignored. By implementing this strategy, we effectively simplify semantic masks and ensure the LLM’s reasoning focuses on relevant categories.
III-B2 Sub-goal-driven top-down map
Considering that the top-down view can better express spatial relationships than the first-person view, we further utilize the depth image $I_{t}^{D}$ to map semantic masks to 3D space, and then convert them into a top-down map. Specifically, by applying the above 2D visual perception process to the RGB image $I_{t}^{R}$, we obtain a segmented image with clearly identified objects and regions. Subsequently, the depth image $I_{t}^{D}$ is converted into a 3D point cloud, where each pixel is mapped to a corresponding 3D point $(X,Y,Z)$ based on its depth value and camera parameters:
@@ III-B2 Sub-goal-driven top-down map
where $(u,v)$ are the pixel coordinates, $(c_{x},c_{y})$ are the camera’s principal point coordinates, and $(f_{x},f_{y})$ are the focal lengths. The semantic labels from the segmented RGB image are mapped to the corresponding 3D point clouds, resulting in a point cloud with semantic information $(X,Y,Z,C_{i})$, where $C_{i}$ are the semantic categories. Then the 3D point cloud is partitioned into discrete voxels, where each voxel aggregates its point clouds as one semantic category using max pooling described in [chaplot2020learning]. Considering that UAVs usually fly above ground objects, for a specific coordinate $(x,y)$, the semantic label of the top one in a column of voxels will be projected into the top-down map:
@@
where $z_{top}$ means the highest $z$ coordinate at location $(x,y)$ and $\text{Voxel}(x,y,z_{top})$ denotes the corresponding semantic label. So, we get a top-down map with semantic information. Notably, since the LLM prioritizes the landmarks in the current sub-goal, if a category contained in the sub-goal appears in the voxel at any $z$ coordinate, this category will be projected into the top-down map first. Thus we can modify Eq. 4 as:
@@
At each step, the position of a UAV will also be recorded as waypoints with special values and projected onto the top-down map.
III-B3 Matrix representation
@@
We found that directly inputting the image-format top-down map to a vision-language model (VLM) often yields poor reasoning results. Alternatively, we process the visual map into a text-based matrix representation, which is input to an LLM to guide the navigation. Specifically, the UAV incrementally collects observations and gradually constructs the top-down map of the surrounding environment. Considering the perception quality of onboard sensors and the effective processing capacity of modern LLMs, we define a local map centered at the UAV’s current position as shown in Fig. 1. Subsequently, we generate a matrix by partitioning the local map into a $20\times 20$ coordinate grid and applying semantic max pooling in each grid, making them finally contain one semantic category. Particularly, we grade all values that fall in each grid in the local map according to frequency, showing only the most frequent features as the final category of this grid. Each category such as building or car is discriminated by different number labels. Furthermore, the grid size $r$ (5m) is also input to the LLM as the metric information (see Table VII for ablation study on the grid size). In this case, the LLM can directly reason about positional relationships among each landmark.
III-C Designing Text Prompt for LLM Planning
Given that LLMs are well-trained in reasoning and planning, we complement their spatial perception and scene understanding capability with formatted prompt and action space. This prompt involves two components. First, the task description is prepared which includes the definition of the environment and the UAV, as well as the format of input and output. We require an LLM to leverage Chain-of-Thought reasoning [wei2022chain] after capturing the observation, thinking step by step in an order of observation-thought-planning-history to predict the next action. Regarding the input instruction, sub-goals are extracted by the LLM, and the landmarks given in the instructions will be updated in the task definition as the scene changes. Second, to improve the quality of prompts, we implemented a cyclic refinement process. This involves designing a new set of instructions for an LLM, including elements such as ‘history action’ and ‘next planning’ corresponding to the STMR module. The task description also emphasizes updating the status of planning, including three states, i.e., todo, in process, and completed, in order to minimize the problem of hallucinations that is common in LLM. The STMR is gradually updated in this iterative process as the UAV explores to improve the accuracy and relevance of the prompts.
Example of LLM Prompt (simplified) [Task Description]You are an embodied unmanned aerial vehicle that navigates in the real world. You need to explore between some places marked and ultimately find the destination to stop.
[Input Format] ‘Instruction’ is a global, step-by-step detailed guidance. ‘History’ is your previously executed actions. ‘Map’ is a top-down coordinate map, using a 20*20 grid M to represent coordinates, heading towards upper of the map. The value of the grid represent different landmarks. You are at [9,9] on the ’Map’ (marked as P). The distance between adjacent numbers is 5 meters in the real world. The numbers in the matrix M represent:
[0:Unexplored 1:your past trajectory 2:road 3:building … ]. ‘Plan’ records previous long-term multi-step planning info that you can refer to now.
[Output Format] ‘Thought’: your thoughts about the task, which may include your comprehension, surroundings, history, and etc. ‘Observation’: the description of the map. ‘Plan’: your updated plan. ‘Action’: your next action.”
Think step by step. First, describe your orientation and the landmark in your left, right and front. Second, according the ’map’, update your multi-step ‘Plan’. Each plan needs to follow a state word (TODO, In Process, Completed). Finally, give the specific “Action” according to the action format: Action: (right, left, lift, down, straight, back), (degree (0-90 degrees)), (distance (between 0-10 meters)).
  A Simplified Input in the Cyclic Process
Instruction
: lift off and head straight across the water to the road…
History: [lift off 5 meters…]
Map: STMR
Plan:
1. Lift off 5 meters. (Completed)
2. Turn 180 degrees. (Completed)
3. Head straight across the water to the road. (Completed)
4. Turn left to align the road. (In Process)
5. (TODO)…
IV Experiment
TABLE I: Comparison on the validation seen of the AerialVLN-S. * denotes we train these methods by running their source codes.
Validation Seen
model NE/m $\downarrow$ SR/$\uparrow$ OSR/$\uparrow$ SDTW/$\uparrow$

### v3

@@ Abstract
Aerial Vision-and-Language Navigation (VLN) is a novel task enabling Unmanned Aerial Vehicles (UAVs) to navigate in outdoor environments through natural language instructions and visual cues. However, it remains challenging due to the complex spatial relationships in aerial scenes. In this paper, we propose a training-free, zero-shot framework for aerial VLN tasks, where the large language model (LLM) is leveraged as the agent for action prediction. Specifically, we develop a novel Semantic-Topo-Metric Representation (STMR) to enhance the spatial reasoning capabilities of LLMs. This is achieved by extracting and projecting instruction-related semantic masks onto a top-down map, which presents spatial and topological information about surrounding landmarks and grows during the navigation process. At each step, a local map centered at the UAV is extracted from the growing top-down map, and transformed into a matrix representation with distance metrics, serving as the text prompt to LLM for action prediction in response to the given instruction. Experiments conducted in real and simulation environments have proved the effectiveness and robustness of our method, achieving absolute success rate improvements of 26.8% and 5.8% over current state-of-the-art methods on simple and complex navigation tasks, respectively. The dataset and code will be released soon.
Introduction
@@
The Aerial Vision-and-Language Navigation (Aerial VLN) (Liu et al. 2023c) emerges as a groundbreaking task. It enables unmanned aerial vehicles (UAVs) to interpret natural language instructions and visual information to navigate in outdoor environments. This technology can eliminate the necessity for manual UAV operation by human pilots, clearly mitigating the barriers to human-UAV interaction and potentially benefitting rescue, search, and delivery tasks.
Figure 1: The pipeline to obtain STMR. (a) The observed RGB image, the corresponding segmented image, and the depth image. (b) Segmented images are projected into the top-down map gradually during the UAV flight, which captures the semantic and topological information of the environment. (c) The top-down map is further transformed into a 20x20 matrix representation with distance metrics for LLM reasoning.
@@
Recently, VLN tasks have been well-developed. Considering the powerful reasoning capability of large language models (LLMs) and vision language models (VLMs) (Ahn et al. 2022; Song et al. 2023; Driess et al. 2023; Liu et al. 2023a, 2024), several VLN methods have started to use LLMs or VLMs as agents to parse instructions and predict actions (Zhou, Hong, and Wu 2024; Shah et al. 2023; Li, Chen, and Lin 2024; Lin et al. 2024). Specifically, LLM-based methods attempt to describe the visual observations with text to enhance the LLM’s scene understanding ability (Zhou, Hong, and Wu 2024; Chen et al. 2024b). VLM-based methods integrate the visual and text data and generates context-aware decisions, including (Zhang et al. 2024c, b; Lee et al. 2024; Sharshar et al. 2025).
@@
Although existing LLM-based and VLM-based methods have made significant progress in indoor or ground-based outdoor environments, they still struggle to effectively encode the large-scale spatial information from the aerial view. As shown in  Fig. 1, the aerial scene can be highly complex, which may lead to the overemphasis of instruction-irrelevant objects or the failure to capture the contextual relationships among different areas. Worse still, existing VLMs suffer from limitations in spatial reasoning when taking raw observations as input (Chen et al. 2024a; Stogiannidis, McDonagh, and Tsaftaris 2025). Thus, there is an urgent need for the development of VLN methods that effectively integrate both semantic information and precise spatial representations to improve adaptability in aerial scenarios.
@@
To overcome these challenges, we propose a zero-shot LLM-based aerial VLN framework that encodes natural language instructions, RGB images and depth images as text prompt, and generates action predictions (e.g., go forward 10 meters) through LLMs directly. Specifically, a Semantic-Topo-Metric Representation (STMR) is designed for LLMs spatial reasoning. Firstly, the instruction-related landmarks are extracted and corresponding semantic masks are generated via visual perception models, i.e., Grounding DINO (Liu et al. 2023b) and Tokenize Anything (Pan et al. 2023). Then, the semantic mask is projected into a top-down map as shown in Fig. 1 (b). This top-down map encompasses both the UAV’s travel trajectory and spatial information, and it grows gradually during the navigation process. To encode the visual information into an LLM-friendly form, we firstly select a fixed-size region centered on the UAV in the top-down map as a local map, then separate the map into equally spaced grids and substitute each grid with a semantic number. As can be seen from Fig. 1 (c), the matrix representation encompasses topological, semantic, and metric information. It is input to an LLM together with historical actions and text instructions to infer the next action. Experiments have shown that the proposed method significantly enhances the capability of spatial reasoning, and achieves absolute success rate improvements of 26.8% and 5.8% over current state-of-the-art methods on simple and complex navigation tasks.
@@
  • To our knowledge, we design the first LLM-based training-free framework for the aerial VLN task, facilitating the development of UAV navigators. Without an extra action planner, the proposed framework allows for easy integration.
@@
Related Work
Vision-Language Navigation (VLN).
@@
VLN aims to enable autonomous agents to navigate complex environments by understanding and executing natural language instructions based on visual context. Early VLN methods use sequence-to-sequence LSTMs to predict low-level actions (Anderson et al. 2018) or high-level actions from panoramas (Fried et al. 2018). Several attention processes have been proposed (Qi et al. 2020; Hong et al. 2020; An et al. 2021) to enhance the process of learning visual textual correspondence. Reinforcement learning is also explored to improve policy learning (Wang et al. 2018; Tan, Yu, and Bansal 2019; Wang, Wu, and Shen 2020). Besides, transformer-based architecture have shown superior performance to long-distance contextual information (Hao et al. 2020; Majumdar et al. 2020). More recent works (Zhou, Hong, and Wu 2024; Chen et al. 2024b) leverage the reasoning and dialogue capabilities of LLMs, achieving great progress. However, most of them operate in limited spaces on the ground. In contrast, aerial VLN remains challenging due to the large-scale and complex environments.
UAV Navigation.
@@
Unmanned Aerial Vehicle (UAV) navigation has seen a surge of interest over the past years. Many earlier works (Blukis et al. 2018a, b) using a combination of supervised and imitation learning for efficient training and low-level velocity prediction to guide UAV’s control actions. LINGUNET (Misra et al. 2018) decomposes instruction execution into two stages, separately using supervised learning for goal prediction and policy gradient for action generation. AerialVLN (Liu et al. 2023c) contributes a much more challenging aerial VLN dataset focusing on outdoor aerial environments, and provides a look-ahead guidance method as the baseline. Recent research (Wang et al. 2024; Sautenkov et al. 2025; Tian et al. 2025) exploiting the powerful visual reasoning capabilities of VLM to address the challenges posed by UAV navigation scenes. Despite the progress, the generalizability and performance of these methods still require improvement.
LLMs for Robot Planning and Interaction.
@@
Most recently, LLMs have demonstrated impressive capabilities in understanding and reasoning. To leverage these capabilities, several promising methods have been proposed for applying LLMs in robotic systems. A few methods involve using LLM-generated rewards optimized in simulation to improve control (Huang et al. 2023; Yu et al. 2023). Others utilize LLM-selected subgoals as an abstraction to enhance policies for navigation (Dorbala et al. 2022; Chen et al. 2023) and manipulation (Cui et al. 2022; Li et al. 2024). Additionally, research has explored the use of LLMs to generate executable code for control and perception primitives (Singh et al. 2023; Liang et al. 2023; Li et al. 2025). Despite their potential, LLMs are still prone to confidently hallucinating outputs, such as referring to objects not observed in the scene (Zeng et al. 2022). In order to alleviate the hallucinating phenomenon, we propose a matrix-based representation containing topological, semantic, and metric information for better prompting the LLMs.
Method
@@
In this paper, we propose a zero-shot framework that leverages Large Language Models (LLMs) for action prediction in aerial VLN tasks. As shown in Fig. 2, our framework consists of three modules. The sub-goal extraction module decomposes language instructions into several sub-goals, facilitating step-by-step reasoning and navigation. The Semantic-Topo-Metric Representation (STMR) module represents the outdoor environment as a matrix containing semantic, topological, and metric information, enhancing the spatial reasoning ability of LLMs. Finally, an LLM planner is designed, which takes sub-goal instructions, STMR, together with historical information and task description as input, and outputs its thoughts and predicted actions.
Problem Formulation
@@
The aerial VLN task is formulated as a free-form language instruction guided navigation. At the beginning of each episode, the initial pose of a UAV is denoted as $P=[x,y,z,\phi,\theta,\psi]$, where $(x,y,z)$ is the UAV’s position and $(\phi,\theta,\psi)$ represents pitch, roll, and yaw of the UAV’s orientation. A natural language instruction $L$ is provided to specify the path that a UAV should follow. To achieve the navigation goal, the UAV considers both the instruction and visual perceptions, and predicts an action from the action space (i.e., right, left, up, down, forward, backward) with corresponding value (e.g., 1-10m) at each time step $t$. Navigation ends when the UAV predicts a ‘Stop’ action or reaches a pre-defined maximum action number. Following (Liu et al. 2023c), the navigation is considered successful if the UAV halts within 20 meters of the target location.
Figure 2: Our method consists of three modules, i.e., Sub-Goal Extraction, Semantic-Topo-Metric Representation, and LLM planner. They are utilized to generate sub-goal instructions, spatial information representations, and UAV navigation actions, respectively.
Semantic-Topo-Metric Representation (STMR)
@@
Previous LLM-based VLN methods use natural language to describe current observations, or a topological graph to model the spatial information of the environment. However, in open scenarios, simple directional words such as “next to” or “aside” are insufficient for describing complex spatial relationships, often leading to ambiguity in LLMs. To address this challenge, we introduce the STMR to enhance the spatial-aware reasoning capability of LLMs. Specifically, STMR incrementally takes an RGB image $I_{t}^{R}$ and a depth map $I_{t}^{D}$ as input from each step, and generates a dynamically updated matrix representation with semantic, topological, and metric information as its output. The details of STMR are presented as follows.
2D Visual Perception.
@@
$\{m^{(t)},\ h^{(t)}\}=\operatorname{TA}\left({I_{t}^{R}},\ \operatorname{GD}\left({I_{t}^{R}},\ \operatorname{LE}({L})\right)\right).\\$ (1)
To improve the robustness of the semantic masks and reduce the misleading of LLM reasoning caused by numerous open-vocabulary categories, we propose a text-matching method to mitigate over-classification. As shown in Fig. 3, our method involves vectorizing the landmarks extracted from the instruction and the captions $h^{(t)}$ generated in each $I_{t}^{R}$. Then, we calculate the cosine similarity between these vectorized landmarks and captions using TF-IDF (Qaiser and Ali 2018). If the similarity score exceeds the threshold $\tau$ (0.8), the landmark is classified as visible in the current view, and other irrelevant masks will be ignored. By implementing this strategy, we effectively simplify semantic masks and ensure the LLM’s reasoning focuses on relevant categories.
Sub-goal-driven Top-down Map.
Considering that a top-down view better expresses spatial relations, we use the depth image to map semantic masks to the 3D space, and then project them into a top-down map. Note that the top-down map grows and updates during the UAV flight to support the navigation process. Specifically, by applying the above 2D visual perception process to the continuously observed RGB images, we obtain segmented images with identified objects and regions. Subsequently, the corresponding depth images are projected to a gradually growing 3D point cloud, where each pixel is mapped to a 3D point $(X,Y,Z)$ based on its depth value and camera parameters:
@@
where $(u,v)$ are the pixel coordinates, $(c_{x},c_{y})$ are the camera’s principal point coordinates, and $(f_{x},f_{y})$ are the focal lengths. The semantic labels from the segmented image are mapped to the corresponding 3D points, resulting in a point cloud with semantic information $(X,Y,Z,C_{i})$, where $C_{i}$ is the semantic category. Then the 3D point cloud is partitioned into discrete voxels, where each voxel aggregates its point clouds as one semantic category using max pooling. For a specific coordinate $(x,y)$, sometimes different categories of objects will appear at different heights. For example, there is ‘vegetation’ or ‘equipment’ on the roof of a ‘building’. Given that UAVs usually fly above the landmarks, the semantic label of the top one in a column of voxels will be projected into the top-down map:
@@
where $z_{top}$ means the highest $z$ coordinate at location $(x,y)$ and $\text{Voxel}(x,y,z_{top})$ denotes the corresponding semantic label. Consequently, we get a top-down map with semantic information. Notably, since the LLM prioritizes the landmarks in the current sub-goal, if a category contained in the sub-goal appears in the voxel at any $z$ coordinate, this category will be projected into the top-down map first. Thus we can modify Eq. 4 as:
@@
Matrix Representation.
@@
We found that directly inputting the image-format top-down map to a vision-language model (VLM) often yields poor reasoning results. Alternatively, we process the visual map into an LLM-friendly text-based matrix representation, which effectively guides the navigation. In order to make LLMs aware of metric information, we define a 100m $\times$ 100m local map centered on the UAV’s current position from the entire top-down map, and divide it into $20\times 20$ coordinate grids with each grid covering a 5m $\times$ 5m area. The grid size is selected according to the size of the most common small landmarks (i.e., cars) to ensure them occupy a single grid, while larger landmarks can be represented by multiple grids. Subsequently, a text-based matrix is generated by applying semantic max pooling in each grid, where the most frequent category is selected as the semantic label. Each category, such as building or car, is identified by different number labels. Furthermore, the grid size $r$ (5m) is also input to the LLM as the metric information. Notably, the local map slides dynamically during the UAV flight, enhancing the LLM to understand the surrounding environment and reason about positional relationships among landmarks.
LLM-based Navigator
To improve the robustness of the LLM-based navigator, the prompt mainly consists of two components, i.e., task definition and STMR-based planner. Firstly, the task definition includes a description of the navigation task, as well as the input and output format. The input format specifies the size of the matrix representation and defines the mapping relationship between the semantic labels and matrix elements. The output format outlines the action space, i.e., (right, left, lift, down, straight, back), along with the maximum moving distance and turning range. Secondly, the STMR-based planner includes a text-based matrix representing contextual observation, sub-goal-based next-step planning, and history actions. The LLMs are required to leverage Chain-of-Thought (Wei et al. 2022) reasoning after capturing the observation, thinking step-by-step in the order of observation-thought-planning-history to predict the next actions. During the navigation process, the status of each sub-goal is updated, consisting of three states, i.e., todo, in process, and completed, to further alleviate the issue of LLM hallucinations. The text-based matrix is progressively updated as the UAV navigates through the environment. The prompt examples can be found in the supplementary material.
Experiment
Dataset and Implementation Details
Dataset.
We conducted experiments using two test sets, i.e., the simple navigation test set collected by us and the complex navigation test set from the AerialVLN-S benchmark (Liu et al. 2023c). Both test sets are derived from Microsoft AirSim plugins and Unreal Engine 4 scenes. Our simple test set consists of 1,000 trajectories that cover a range of common navigation scenarios, e.g., public transportation and urban environments. Each trajectory spans up to 100 meters and contains 1 or 2 landmarks as navigation references. This test set will also be released with the code. AerialVLN-S replicates real-world urban environments with over 870 object categories across diverse scenarios, including downtown cities, factories, parks, and villages. The flight trajectories are collected by AOPA-certified UAV pilots and curated by experts to ensure realism and navigational fidelity. For the AerialVLN-S benchmark, the trajectory is much longer (326.9 meters on average) with more landmarks and more complex instructions (83 words per instruction on average), causing an extremely challenging task. Besides, to verify the robustness and generalizability of the proposed method, we conduct real-world experiments in 10 outdoor scenes.
Evaluation Metrics.
To validate the effectiveness of our method, we utilize a comprehensive set of evaluation metrics following (Liu et al. 2023c). We focus on several key aspects, i.e., Navigation Error (NE), quantifying the distance between the UAV’s stopping point and the actual destination; Success Rate (SR), measuring the proportion of navigations that successfully reach the destination within a 20-meter threshold; Oracle Success Rate (OSR), an idealized measure considering any point on the predicted trajectory that comes within 20 meters of the destination as a success; Normalised Dynamic Time Warping (SDTW), taking into account both the navigation success rate and the similarity between the predicted trajectory and ground truth.
Implementation Details.
Our framework is implemented in both a simulator and a real environment. In the simulation environments of AirSim and UE4, we evaluated the methods using a PC equipped with an Intel i9 12th-generation CPU and Nvidia RTX 4090 GPUs. For the real world environment, we test on a Q250 airframe, carrying an Intel RealSense D435i depth camera, and an NVIDIA Jetson Xavier NX running Ubuntu 18.04 as the onboard computer. This setup supports the construction of top-down maps and flight control. The deployment of large-scale models and communication with the onboard computer are managed by a mobile ground station equipped with high-performance computing capabilities (matching the simulation setup). For LLM reasoning, we utilize the online API of GPT-4o, employing the default parameters.
Figure 4: Demonstration of the STMR in spatial reasoning.
model NE/m $\downarrow$ SR/% $\uparrow$ OSR/% $\uparrow$ SDTW/% $\uparrow$
Random 266.7 2.0 5.9 0.2
Action Sampling 259.9 5.1 8.0 0.3
Seq2Seq (Anderson et al. 2018) 130.1 25.5 57.1 10.6
CMA (Hong et al. 2022) 113.6 29.8 73.2 10.5
AerialVLN (Liu et al. 2023c) 116.4 37.7 71.7 12.7
Navid (Zhang et al. 2024a) 99.5 44.1 79.9 16.6
OpenUAV (Wang et al. 2024) 108.8 49.9 68.0 14.9
Ours (training-free) 53.9 76.7 90.3 18.0
Table 1: Overall performance comparisons on simple tasks.
AerialVLN-S
model NE/m $\downarrow$ SR/% $\uparrow$ OSR/% $\uparrow$ SDTW/% $\uparrow$